# Spectral properties from an efficient analytical representation of the $GW$ self-energy within a multipole approximation

### v1

@@ Abstract
We propose an efficient analytical representation of the frequency-dependent $GW$ self-energy $\Sigma$ via a multipole approximation (MPA-$\Sigma$). The multipole-Padé model for the self-energy is interpolated from a small set of numerical evaluations of $\Sigma$ in the complex frequency plane, similarly to the previously multipole representation developed for the screened Coulomb interaction (MPA-$W$) [Phys. Rev. B 104, 115157 (2021)]. We show that, likewise MPA-$W$, an appropriate choice of frequency sampling in MPA-$\Sigma$ is critical to guarantee computational efficiency and high accuracy. The combined MPA-$W$ and MPA-$\Sigma$ scheme considerably reduces the cost of full-frequency self-energy calculations, especially for spectral band structures over a wide energy range. Crucially, MPA-$\Sigma$ enables a multipole representation for the interacting Green’s function $G$ (MPA-$G$), providing a straightforward evaluation of all the spectral properties, and a more general way to define the renormalization factor $Z$. We validate the MPA-$\Sigma$ and MPA-$G$ approaches for diverse systems: bulk Si, Na and Cu, monolayer MoS2, the NaCl ion-pair and the F2 molecule. Moreover, we introduce toy MPA-$\Sigma$/$G$ models to examine the quasiparticle picture in different regimens of weak and strong correlation. With these models, we expose limitations in defining $Z$ from the local derivative of $\Sigma$.
@@ I Introduction
Within the framework of first-principle methods in condensed matter physics, mean field approaches such as density functional theory (DFT) in the Kohn-Sham (KS) approximation give accurate ground state properties and have been immensely useful for understanding the electronic structure of materials. However, it fails to reliably provide accurate band structures, as this requires including many-body effects beyond the DFT level. The description of electron addition or removal energies and related excited state properties is usually treated with methods such as the $GW$ approximation, based on the Green’s function formalism [1, 2, 3, 4, 5, 6].
@@
In common $GW$ implementations, the Green’s function $G$ and the screened Coulomb potential $W$ are constructed perturbatively. Starting from DFT, the KS quasiparticle (QP) energies are corrected by an exchange-correlation self-energy $\Sigma$ computed at the $GW$ level. This correction can be iterated within a self-consistent $GW$ approach, or done in a computationally cheaper one-shot $G_{0}W_{0}$ procedure. Since the imaginary part of $G$ is closely related to the spectral function obtained from photoemission experiments [7, 8], a dynamical self-energy $\Sigma(\omega)$ can account for many-body features, such as finite QP lifetimes and satellite structures [2, 9, 10, 11, 12, 13, 5, 14].
@@
$G_{0}W_{0}$ is the state-of-the-art ab initio method for the description of angle-resolved photoemission and inverse photoemission spectroscopy measurements, giving generally a very accurate agreement with experiment (see e.g. Refs. [15, 16, 17, 18]). More accurate spectral functions can be obtained with self-consistent approaches, including cumulant expansions of $\Sigma$ and vertex corrections (see e.g. Refs. [11, 4, 19, 20, 21, 5, 14]).
@@
The $GW$ self-energy $\Sigma(\omega)$ is given by a frequency convolution of $G(\omega)$ and $W(\omega)$. This convolution can be evaluated with different full-frequency (FF) methods, based on numerical integrations along the frequency real axis [22, 23, 17, 24], or through an integration in the complex frequency plane using contour deformation and analytic continuation techniques [25, 26, 27, 28, 29, 30]. Such numerical FF evaluations tend to be computationally expensive. A less costly alternative is to integrate $\Sigma$ analytically after approximating the frequency dependence of $W$ (or the dielectric function) with simple models, such as the plasmon pole approximation (PPA) [31, 32, 33, 34, 35], which in many cases has limited accuracy. A higher accuracy can be obtained with multipole models and Padé approximants [36, 29, 37, 38], including the recently developed MPA-$W$ method [39, 40, 41].
@@
Analytical models of the dielectric response are also widely used in the study of optical and electronic properties of materials [42, 43, 44]. They have been used in the study of, e.g., optical excitations [45, 46], electron energy loss [47, 48], and X-ray absorption spectra [49, 50]. Simple models have also been used to account for dynamical effects arising from electron-hole interactions in doped systems [51, 52], or in the ab initio description of the plasmon-phonon hybridization in doped semiconductors [53]. Less common is the use of models for $\Sigma$ [54, 55, 56, 57], the interacting $G$ [57, 58, 59], or for total energy calculations [60, 57, 58, 59, 61], which usually aims mainly at improving the computational efficiency of such calculations.
@@
In this work, we present an efficient multipole approximation for the self-energy (MPA-$\Sigma$). This method yields simple analytical representations of all the $GW$ operators, including a multipole-Padé representation of the Green’s function (MPA-$G$). Moreover, the combination of MPA-$\Sigma$ with the previous MPA-$W$ method considerably reduces the cost of evaluating $\Sigma$ and $G$ in its full-frequency domain. Although we limited the study to the $G_{0}W_{0}$ approximation, its extension to higher levels of theory, e.g. self-consistent $GW$, and the inclusion of vertex corrections or cumulant expansions, is straightforward. Within the same framework, we present MPA toy models that provide new insight into the QP picture and the renormalization factor $Z$.
@@
The paper is organized as follows: in the methods section (Sec. II) we summarize the general $GW$ equations (Sec. II.1), MPA-$W$ (Sec. II.2), the new MPA-$\Sigma$ (Sec. II.3) and MPA-$G$ (Sec. II.4) approaches, and provide computational details (Sec. II.5). In the results section (Sec. III) we benchmark MPA-$\Sigma$ and MPA-$G$ on different prototypical materials (Sec. III.1), build spectral band structures (Sec. III.2), and finally analyze the QP picture using toy MPA models (Sec. III.3). Sec. IV holds our conclusions.
@@ II.1 Quasiparticle $GW$ equations
$G_{0}(z)=\sum_{m}P_{m}^{\text{KS}}\left[\frac{f_{m}^{\text{KS}}}{z^{*}-\epsilon_{m}^{\text{KS}}}+\frac{(1-f_{m}^{\text{KS}})}{z-\epsilon_{m}^{\text{KS}}}\right],$ (1)
where the sum runs over the KS states, $m$, with the projectors $P_{m}^{\text{KS}}=|\psi_{m}^{\text{KS}}\rangle\langle\psi_{m}^{\text{KS}}|$, KS energies $\epsilon_{m}^{\text{KS}}$ and occupation numbers $f_{m}^{\text{KS}}\in[0,1]$. The complex frequency is given by $z\equiv\omega+i\varpi$ (see notation in Table 1), with $\varpi>0$ to ensure the time ordering.
@@
$G_{0n{\mathbf{k}}}(z)\equiv$ $\langle\psi_{n{\mathbf{k}}}^{\text{KS}}|G_{0}(z)|\psi_{n{\mathbf{k}}}^{\text{KS}}\rangle=$ (2)
$\frac{f_{m}^{\text{KS}}}{z^{*}-\epsilon^{\text{KS}}_{n{\mathbf{k}}}}+\frac{(1-f_{m}^{\text{KS}})}{z-\epsilon^{\text{KS}}_{n{\mathbf{k}}}},$
where the spectral function $\mathrm{Im}~{}[G_{0n{\mathbf{k}}}]$ is a Dirac delta function centered on $\epsilon^{\text{KS}}_{n{\mathbf{k}}}$. The interacting Green’s function is given by the Dyson equation for this operator:
$G_{n{\mathbf{k}}}^{-1}(z)=G_{0n{\mathbf{k}}}^{-1}(z)-\Sigma_{n{\mathbf{k}}}(z),$ (3)
@@
The QP energies correspond to the poles of $G$ in the $\varpi\to 0^{+}$ limit, which are determined by solving the QP equation:
@@
The frequency dependence of $G$ has a structure typically dominated by a well-defined main peak, the QP pole [8], and satellite structures at larger energies [63, 4, 5]. Like the QP pole, the satellites are also formal solutions of Eq. (5). They arise from the plasmonic structures of $\Sigma$ and give rise to replicas of the QP band structure [4, 5]. In the so-called QP picture, satellites are disregarded and only energies around the QP pole are considered. As such, the QP picture resembles the independent particle picture, but with the KS energies corrected by the real part of $\Sigma$, while the finite imaginary part accounts for the broadening of the QP pole, according to its lifetime.
@@
$G(z)$ $\epsilon_{p}=\varepsilon_{p}+i\eta_{p}$ $P_{p}$
@@
Table 1: List of complex quantities relevant for this work, and definition of the used notation.
@@
$\epsilon_{n{\mathbf{k}}}\approx\epsilon^{\text{lin}}_{n{\mathbf{k}}}\equiv\epsilon_{n{\mathbf{k}}}^{\text{KS}}+Z_{n{\mathbf{k}}}\langle\psi_{n{\mathbf{k}}}^{\text{KS}}|\Sigma(\epsilon_{n{\mathbf{k}}}^{\text{KS}})-v_{xc}^{\text{KS}}|\psi_{n{\mathbf{k}}}^{\text{KS}}\rangle,$ (6)
with the corresponding renormalization factor, $Z_{n{\mathbf{k}}}$, given by
$Z_{n{\mathbf{k}}}=\left[1-\langle\psi_{n{\mathbf{k}}}^{\text{KS}}|\frac{\partial\Sigma(\omega)}{\partial\omega}\bigg{|}_{\omega=\epsilon_{n{\mathbf{k}}}^{\text{KS}}}|\psi_{n{\mathbf{k}}}^{\text{KS}}\rangle\right]^{-1},$ (7)
which approximates the spectral weight of the QP pole in the limit of weak correlation. Since the satellite structures have a non-vanishing weight, $Z_{n{\mathbf{k}}}$ typically has a very small imaginary part and a real part ranging from 0.5 to 1. This interval gives the validity range of the QP picture [63, 5, 64]. When computed in a consistent way, the spectral weights of the QP pole and the satellites sum exactly to one, since they comply with the sum rule for the number of particles/holes [65]:
@@ II.2 MPA for the screening interaction
The screened Coulomb potential can be separated in a static bare Coulomb and a correlation term: $W(\omega)=v+W_{\text{c}}(\omega)$. As detailed in Ref. [39], the frequency dependence of each matrix element $W^{\text{c}}_{{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}$ can be described by a multipole model with a small number of complex poles for each transferred momentum, ${\mathbf{q}}\equiv{\mathbf{k}}^{\prime}-{\mathbf{k}}$, and reciprocal lattice vectors ${\mathbf{G}}{\mathbf{G}}^{\prime}$. $W$ is then given by (indexes ${\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}$ omitted)
$W^{\text{MPA}}(z)=v+\sum_{p=1}^{n_{W}}R_{p}\left[\frac{1}{z-\Omega_{p}}-\frac{1}{z+\Omega_{p}}\right].$ (9)
Here $\Omega_{p}$ are the MPA poles and $R_{p}$ their residues, and $n_{W}$ the number of poles. The time ordering of $W$ implies that $\mathrm{Re}~{}[\Omega_{p}]\times\mathrm{Im}~{}[\Omega_{p}]<0$. Such poles represent effective plasmon-like quasiparticles emerging from a large set of single-particle transitions from valence to conduction states [36, 40].
For each $W^{\text{c}}_{{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}$ matrix element, all poles and residues are obtained through a non-linear interpolation of values numerically evaluated in a conveniently selected set of complex frequencies $\{z_{j},j=1,\ldots,2n_{W}\}$. We use a frequency sampling along two lines parallel to the real axis (double parallel sampling), one closer (typically with an imaginary part of $0.1$ Ha) and one further away (at $1$ Ha), with an inhomogeneous distribution along the positive real axis (see Eq. (10) of Ref. [40]). The double parallel sampling, in particular the line of points with the largest imaginary part, reduces the noise resulting from the coarse Brillouin zone sampling of $W$. Moreover, the inhomogeneous distributions along the real axis are denser closer to the origin and limit the number of poles needed [39, 40].
The frequency integral in the $G_{0}W_{0}$ self-energy of Eq. (4) can then be integrated analytically, resulting in
$\Sigma^{\text{MPA-}W}(z)=\Sigma_{\text{x}}+\\ \sum_{m}\sum_{p}^{n_{W}}P_{m}^{\text{KS}}R_{p}\left[\frac{f_{m}^{\text{KS}}}{z^{*}-\epsilon_{m}^{\text{KS}}+\Omega_{p}}+\frac{(1-f_{m}^{\text{KS}})}{z-\epsilon_{m}^{\text{KS}}-\Omega_{p}}\right].$ (10)
Note that the time ordering of $W$ carries over to the time ordering of $\Sigma$. The renormalization factor can also be computed analytically as
@@
$(Z^{\text{MPA-}W})^{-1}=-\sum_{m}\sum_{p}^{n_{W}}P_{m}^{\text{KS}}R_{p}\times\\ \left[\frac{f_{m}^{\text{KS}}}{(\epsilon_{m}^{\text{KS}}-\Omega_{p})^{2}}+\frac{(1-f_{m}^{\text{KS}})}{(\epsilon_{m}^{\text{KS}}+\Omega_{p})^{2}}\right].$ (11)
By increasing $n_{W}$ from $1$ to typically about $10$, the MPA-$W$ method goes from a standard single-pole PPA, to an effective full-frequency approach [39]. The MPA-$W$ method is currently implemented in yambo [66, 67] and gpaw [68].

### v2

@@ Abstract
We propose an efficient analytical representation of the frequency-dependent $GW$ self-energy $\Sigma$ via a multipole approximation (MPA-$\Sigma$). The MPA self-energy model is interpolated from a small set of numerical evaluations of $\Sigma$ in the complex frequency plane, similar to the MPA interpolation developed for the screened Coulomb interaction (MPA-$W$) [D. A. Leon et al., Phys. Rev. B 104, 115157 (2021)]. Crucially, MPA-$\Sigma$ enables a multipole representation for the interacting Green’s function $G$ (MPA-$G$), and in turn, access to all the spectral properties, including quasiparticle energies (QP) and renormalization factors beyond the linearized QP equation. We validate the MPA-$\Sigma$ and MPA-$G$ approaches for a diverse set of systems: bulk Si, Na and Cu, monolayer MoS2, the NaCl ion-pair, and the F2 molecule. We show that, just as for MPA-$W$, an appropriate choice of frequency sampling in MPA-$\Sigma$ is critical to guarantee computational efficiency and high accuracy. Moreover, the combined MPA-$W$ and MPA-$\Sigma$ scheme considerably reduces the cost of full-frequency self-energy calculations, especially for spectral band structures over a wide energy range.
@@ I Introduction
In condensed matter physics, first principle methods such as density functional theory (DFT) in the Kohn-Sham (KS) approximation provide accurate ground-state properties and have been immensely useful for understanding the electronic structure of materials. However, they fail to reliably provide accurate band structures, which requires including many-body effects beyond the mean field DFT level. The description of electron addition or removal energies and related excited-state properties is usually treated with methods such as the $GW$ approximation, based on the Green’s function formalism [1, 2, 3, 4, 5, 6].
@@
In common $GW$ implementations, the Green’s function $G$ and the screened Coulomb potential $W$ are constructed perturbatively. Starting from DFT, the KS quasiparticle (QP) energies are corrected by an exchange-correlation self-energy $\Sigma$. This correction can be obtained iteratively within a self-consistent $GW$ approach, or done in a computationally cheaper one-shot $G_{0}W_{0}$ procedure. Since the imaginary part of $G$ is closely related to the spectral function obtained from photoemission experiments [7, 8], a dynamical self-energy $\Sigma(\omega)$ can account for many-body features, such as finite QP lifetimes and satellite structures [9, 10, 2, 11, 12, 13, 14, 15, 5, 16].
@@
$G_{0}W_{0}$ is the state-of-the-art ab initio method for the description of angle-resolved photoemission and inverse photoemission spectroscopy measurements, giving generally a very accurate agreement with experiment (see, e.g., Refs. [17, 18, 19, 20]). More accurate spectral functions can be obtained with self-consistent approaches, including cumulant expansions of $\Sigma$ and vertex corrections (see, e.g., Refs. [13, 4, 21, 22, 23, 5, 16]).
@@
The $GW$ self-energy $\Sigma(\omega)$ is given by a frequency convolution of $G(\omega)$ and $W(\omega)$. This convolution can be evaluated with different full-frequency (FF) methods, based on numerical integrations along the frequency real axis [24, 25, 19, 26], or through an integration in the complex frequency plane using contour deformation and analytic continuation techniques [27, 28, 29, 30, 31, 32]. Such numerical FF evaluations tend to be computationally expensive. A less costly alternative is to represent $W$ (or the dielectric function) with a simple model, such as in the plasmon pole approximation (PPA), that allows for an analytical integration of the frequency convolution in $\Sigma$ [33, 34, 35, 36, 37], but in many cases this has limited accuracy. Higher accuracy can be obtained with multipole models and Padé approximants [38, 31, 39, 40], including the recently developed MPA-$W$ method [41, 42, 43].
@@
Analytical models of the dielectric response are also widely used in the study of optical and electronic properties of materials [44, 45, 46]. They have been used in the study of, e.g., optical excitations [47, 48], electron energy loss [49, 50], and x-ray absorption spectra [51, 52]. Simple models have also been used to account for dynamical effects arising from electron-hole interactions in doped systems [53, 54], or in the ab initio description of plasmon-phonon hybridization in doped semiconductors [55]. Less common is the use of models for $\Sigma$ [56, 57, 58, 59], the interacting $G$ [59, 60, 61, 62], or for total energy calculations [63, 59, 60, 61, 62, 64], and usually these approaches mainly aim at improving the computational efficiency of such calculations.
@@
In this work, we present an efficient multipole approximation for the self-energy (MPA-$\Sigma$). This method yields simple analytical representations of all $GW$ operators, including a multipole-Padé representation of the Green’s function (MPA-$G$). Moreover, the combination of MPA-$\Sigma$ with the previous MPA-$W$ method considerably reduces the cost of evaluating $\Sigma$ and $G$ in its full-frequency domain. Although this study is limited to the $G_{0}W_{0}$ approximation, extending it to higher levels of theory, e.g., self-consistent $GW$, and the inclusion of vertex corrections or cumulant expansions, is straightforward.
@@
The paper is organized as follows: In the methods section (Sec. II) we summarize the general $GW$ equations (Sec. II.1), and the previously introduced MPA-$W$ method (Sec. II.2). Thereafter, we present the new MPA-$\Sigma$ (Sec. II.3) and MPA-$G$ (Sec. II.4) approaches, and provide computational details of the $GW$ calculations (Sec. II.5). In the results section (Sec. III) we benchmark MPA-$\Sigma$ and MPA-$G$ on different prototypical materials (Sec. III.1), and build spectral band structures (Sec. III.2). Last Sec. IV holds our conclusions. In addition, Appendix A provides an analysis of the QP particle and the renormalization factor in terms of two MPA toy models, while Appendix B presents numerical details for interpolating spectral functions in momentum space.
@@ II.1 Quasiparticle $GW$ equations
$G_{0}(z)=\sum_{m}\rho_{m}^{\text{KS}}\left[\frac{f_{m}^{\text{KS}}}{z-\varepsilon_{m}^{\text{KS}}-i0^{+}}+\frac{1-f_{m}^{\text{KS}}}{z-\varepsilon_{m}^{\text{KS}}+i0^{+}}\right],$ (1)
where the sum runs over the KS states, $m$, with the projectors $\rho_{m}^{\text{KS}}=|\psi_{m}^{\text{KS}}\rangle\langle\psi_{m}^{\text{KS}}|$, KS energies $\varepsilon_{m}^{\text{KS}}$, and occupation numbers $f_{m}^{\text{KS}}\in[0,1]$. The complex frequency is given by $z\equiv\omega+i\varpi$, which is evaluated in the first and third quadrants ($\omega\varpi>0$), opposite to the pole position according to the time ordering (see notation in Table 1).
@@
$G_{0n{\mathbf{k}}}(z)$ $\equiv\matrixelement{\psi_{n{\mathbf{k}}}^{\text{KS}}}{G_{0}(z)}{\psi_{n{\mathbf{k}}}^{\text{KS}}}$ (2)
$=\frac{f_{n{\mathbf{k}}}^{\text{KS}}}{z-\varepsilon^{\text{KS}}_{n{\mathbf{k}}}-i0^{+}}+\frac{1-f_{n{\mathbf{k}}}^{\text{KS}}}{z-\varepsilon^{\text{KS}}_{n{\mathbf{k}}}+i0^{+}},$
where the spectral function $\mathrm{Im}[G_{0n{\mathbf{k}}}]$ is a Dirac delta function centered on $\epsilon^{\text{KS}}_{n{\mathbf{k}}}$. The interacting Green’s function is given by the Dyson equation for this operator, in which the DFT exchange and correlation potential is subtracted from the self-energy:
$G_{n{\mathbf{k}}}^{-1}(z)=G_{0n{\mathbf{k}}}^{-1}(z)-\Sigma_{n{\mathbf{k}}}(z)+v_{xc}^{\text{KS}},$ (3)
@@
The QP energies correspond to the poles of $G$, which are determined by solving the QP equation:
@@
The frequency dependence of $G$ has a structure typically dominated by a well-defined main peak, the QP pole [8], and satellite structures at larger energies [66, 4, 5]. Like the QP pole, the satellites are also formal solutions of Eq. (5). They arise from many-body excitations accounted for in $\Sigma$, such as the plasmonic structures in $GW$, and give rise to replicas of the QP band structure [4, 5]. In the so-called QP picture, satellites are disregarded and only energies around the QP pole are considered. As such, the QP picture resembles the independent particle picture, but with the KS energies corrected by the real part of $\Sigma$, while the finite imaginary part accounts for the broadening of the QP pole, according to its lifetime.
@@
$G_{0}(z)$ $\epsilon^{\text{KS}}=\varepsilon^{\text{KS}}\pm i0^{+}$ $1$
$G(z)$ $\epsilon_{p}=\varepsilon_{p}+i\eta_{p}$ $Z_{p}$
@@
Table 1: List of complex quantities relevant for this work, and definition of the used notation. In the case of $G_{0}$, each state is represented by a single pole with a vanishing imaginary part whose sign follows the time ordering. The residue of such pole carries all the spectral weight.
@@
$\epsilon_{n{\mathbf{k}}}\approx\epsilon^{\text{lin}}_{n{\mathbf{k}}}\equiv\epsilon_{n{\mathbf{k}}}^{\text{KS}}+Z^{\text{lin}}_{n{\mathbf{k}}}\langle\psi_{n{\mathbf{k}}}^{\text{KS}}|\Sigma(\epsilon_{n{\mathbf{k}}}^{\text{KS}})-v_{xc}^{\text{KS}}|\psi_{n{\mathbf{k}}}^{\text{KS}}\rangle,$ (6)
with the corresponding linearized renormalization factor, $Z^{\text{lin}}_{n{\mathbf{k}}}$, given by
$Z^{\text{lin}}_{n{\mathbf{k}}}=\left[1-\langle\psi_{n{\mathbf{k}}}^{\text{KS}}|\frac{\partial\Sigma(z)}{\partial z}\bigg{|}_{z=\epsilon_{n{\mathbf{k}}}^{\text{KS}}}|\psi_{n{\mathbf{k}}}^{\text{KS}}\rangle\right]^{-1},$ (7)
which approximates the spectral weight of the QP pole, $Z_{n{\mathbf{k}}}$, based on the assumption that the QP correction ($\epsilon_{n{\mathbf{k}}}-\epsilon^{\text{KS}}_{n{\mathbf{k}}}$) is small. Since the satellite structures have a non-vanishing weight, $Z_{n{\mathbf{k}}}$ usually has a very small imaginary part and a real part ranging from 0.5 to 1. This interval is taken as a typical validity range of the QP picture [66, 5, 67], while strong correlation effects can lead to situations where the spectral weight is concentrated in the satellites, like in Mott insulators (see, e.g., Refs. [68, 4]). When computed in a consistent way, the spectral weights of the QP pole and the satellites sum exactly to one, since they comply with the sum rule for the number of particles and holes [69]:
@@ II.2 MPA for the screening interaction
The screened Coulomb potential can be separated in a static bare Coulomb and a correlation term: $W(\omega)=v+W_{\text{c}}(\omega)$. As detailed in Ref. [41], the frequency dependence of each matrix element $W^{\text{c}}_{{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}$ can be described by a multipole model with a small number of complex poles for each transferred momentum, ${\mathbf{q}}$, and reciprocal lattice vectors, ${\mathbf{G}}{\mathbf{G}}^{\prime}$. $W$ is then given by:
@@
$W^{\text{MPA}}_{{\mathbf{G}}{\mathbf{G}}^{\prime}}({\mathbf{q}},z)=v_{{\mathbf{G}}{\mathbf{G}}^{\prime}}({\mathbf{q}})+\sum_{p=1}^{n_{W}}R_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}\\ \times\left[\frac{1}{z-\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}}-\frac{1}{z+\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}}\right],$ (9)
where $\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}$ are the MPA poles and $R_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}$ their residues, and $n_{W}$, the number of poles. The time ordering of $W$ implies that $\mathrm{Re}[\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}]\times\mathrm{Im}[\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}]<0$. Such poles represent effective plasmon-like quasiparticles emerging from a large set of single-particle transitions from valence to conduction states [38, 42].
For each $W^{\text{c}}_{{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}$ matrix element, all poles and residues are obtained through a non-linear interpolation of values numerically evaluated in a conveniently selected set of complex frequencies $\{z_{i},i=1,\ldots,2n_{W}\}$. We use a frequency sampling along two lines parallel to the real axis (double-parallel sampling), typically along $\mathrm{Im}~{}z=0.1$ and $\mathrm{Im}~{}z=1$ Ha, respectively, with $\mathrm{Re}~{}z_{i}$ distributed inhomogeneously. The double-parallel sampling, in particular the line of points with the largest imaginary part, reduces the noise resulting from the coarse Brillouin zone sampling of $W$. The inhomogeneous sampling distributions along the real axis are denser closer to the origin, following Eq. (10) of Ref. [42], which limits the number of poles needed [41, 42]. We use two types of distributions, a linear and a quadratic semi-homogeneous partition, depending on the given system. The frequency range is also specific for each system, since it must encompass the main structures of $W$. The classical plasmon energy, or the maximum single-particle transition from the valence to the conduction bands can be used as a reference energy scale in setting the sampling. More practical details, including a measure of the representability error, can be found in Refs. [41, 42].
With such an MPA representation, the frequency integral in the $G_{0}W_{0}$ self-energy of Eq. (4) can be solved analytically. In a plane-wave basis set, it results in
$\Sigma_{n{\mathbf{k}}}^{\text{MPA-}W}(z)=\Sigma_{n{\mathbf{k}}}^{\text{x}}+\sum\limits_{m}\sum\limits_{{\mathbf{G}}{\mathbf{G}}^{\prime}}\sum\limits_{p=1}^{n_{W}}\int\frac{d{\mathbf{q}}}{(2\pi)^{3}}S^{nm}_{p{\mathbf{G}}{\mathbf{G}}^{\prime}}({\mathbf{k}},{\mathbf{q}})\\ \times\left[\frac{f_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}}{z-\epsilon_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}+\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}}+\frac{1-f_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}}{z-\epsilon_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}-\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}}\right],$ (10)
where
$S^{nm}_{p{\mathbf{G}}{\mathbf{G}}^{\prime}}({\mathbf{k}},{\mathbf{q}})$ $\equiv-2\rho_{nm}^{\text{KS}}({\mathbf{k}},{\mathbf{q}},{\mathbf{G}})R_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}{\rho_{nm}^{\text{KS}}}^{*}({\mathbf{k}},{\mathbf{q}},{\mathbf{G}}^{\prime})$ (11)
$\rho_{nm}^{\text{KS}}({\mathbf{k}},{\mathbf{q}},{\mathbf{G}})$ $\equiv\matrixelement{n\mathbf{k}}{e^{i(\mathbf{q}+\mathbf{G})\cdot\mathbf{r}}}{m\mathbf{k-q}}.$
Note that the time ordering of $W$ carries over to the time ordering of $\Sigma$, and since $\mathrm{Im}[\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}}]$ is finite the vanishing imaginary part of the $\epsilon_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}$ poles of $G_{0}$ can be disregarded. The derivative of $\Sigma^{\text{MPA-}W}(z)$ and therefore the linearized renormalization factor [see Eq. (7)] can also be computed analytically as
$\frac{\partial\Sigma_{n{\mathbf{k}}}^{\text{MPA-}W}(z)}{\partial z}=-\sum\limits_{m}\sum\limits_{{\mathbf{G}}{\mathbf{G}}^{\prime}}\sum\limits_{p=1}^{n_{W}}\int\frac{d{\mathbf{q}}}{(2\pi)^{3}}S^{nm}_{p{\mathbf{G}}{\mathbf{G}}^{\prime}}({\mathbf{k}},{\mathbf{q}})\\ \times\left[\frac{f_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}}{(z-\epsilon_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}+\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}})^{2}}+\frac{1-f_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}}{(z-\epsilon_{m{\mathbf{k}}-{\mathbf{q}}}^{\text{KS}}-\Omega_{p{\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}})^{2}}\right].$ (12)
The case $n_{W}=1$ is analogous to the PPA approach, which only uses one or two frequency evaluations of $W$. On the other hand, FF approaches on the real-axis can require as much as 1000 frequency points. By increasing $n_{W}$, MPA-$W$ typically provides an accuracy similar to FF with about 10 poles, interpolated from 20 frequency points. Thus, MPA-$W$ can be viewed as an effective FF approach requiring around 50 times fewer $W$ evaluations, with the corresponding savings in memory allocation [41]. The MPA-$W$ method is currently implemented in yambo [70, 71] and gpaw [72].